# Understanding Knowledge Hijack Mechanism in In-context Learning through Associative Memory

### v1

@@ 4.1 Oversight of in-context knowledge
For vocabulary $v\neq v_{1},v_{2}$, the logit $\xi_{v}$ is determined only by $\log{\pi_{b}(v\mid q)}$ and this means that the candidate for the next token is $\operatorname*{arg\,max}_{v\in\mathcal{V}\setminus\{v_{1},v_{2}\}}\pi_{b}(v\mid q)$. On the other hand, The logits for $v_{1}$ and $v_{2}$ have the other term, which makes them other candidates for the next token. Overall, Corollary 1 states that the model either uses global bigram knowledge or in-context knowledge, depending on the bigram conditionals $\pi_{b}(\cdot\mid q)$ and the frequency of token patterns $qv_{1}$ and $qv_{2}$ in the context.
@@ 5 Experiments
Henceforth, we denote by $\operatorname{TF}_{\text{ape}},\operatorname{TF}_{\text{rpe}}$ and $\operatorname{TF}_{\text{orc}}$ two-layer transformer with APE, RPE, and oracle two-layer transformer parameterized by $\tau_{1},\tau_{2}$ as in Def. 3, respectively. Note that we construct $\operatorname{TF}_{\text{orc}}$ by setting appropriate weights for each matrix in the transformer. We used the two-layer transformer architecture $\operatorname{TF}_{\text{ape}}$ constructed in Sec. 3.1.
@@ 5.1 Oversight of in-context knowledge
Figure 2: Comparison of a two-layer transformer with APE and RPE in capturing previous token information. Even when trained with sequences of length $256$, the attention block in the first layer of $\operatorname{TF}_{\text{APE}}$ fails to attend to the previous token at positions $t>128$, while $\operatorname{TF}_{\text{RPE}}$ attends to previous tokens regardless of the positions.
@@
We investigated whether the training of transformers directs attention towards the previous token. Specifically, $\operatorname{TF}_{\text{rpe}}$ and $\operatorname{TF}_{\text{ape}}$ were trained using inputs of length $256$ generated according to the sequence generation rule. Details regarding the hyperparameters and sequence generation process are provided in Appendix D.
For an associative memory $\sum_{(i,j)\in M}u_{i}v_{j}^{\top}$, we employed the memory recall metric [12, 18, 4]:
@@
For example, $\operatorname{TF}_{\text{rpe}}$ calculates the attention with $W_{K}^{1}=\sum_{k\in Q}w_{E}(k)r_{-1}^{\top}$, and $W_{Q}^{1}=I$ in the first layer attention block. We want each vocabulary $w_{E}(k)$ to be paired with the information of the previous position $r_{-1}$, not other positions $r_{-i}$ where $i\neq 1$. Hence, $R(W_{K}^{1})$ is computed as
$\frac{1}{\|Q\|}\sum_{k\in Q}\mathbf{1}\{\operatorname*{arg\,max}_{i^{\prime}}r_{-i^{\prime}}^{\top}W_{K}^{1}w_{E}(k)=-1\}.$
For $\operatorname{TF}_{\text{ape}}$, we similarly have
@@
where $T$ is the maximum input length for $\operatorname{TF}_{\text{ape}}$.
As we can see from Fig. 2, over $90\%$ of desired associations are stored in $W_{K}^{1}$ for $\operatorname{TF}_{\text{rpe}}$, while $\operatorname{TF}_{\text{ape}}$ fails to remember most of the associations at positions $256128<t<$ 256 even though the model was trained with sequences of length $256$.
@@ Discussion
Figure 3: Ratio of predicting $B_{1}$ and $B_{2}$ given the prompt "$A\,B_{1}\,,\dots,\,A\,B_{1}\,,\,A\,B_{2}\,,\dots,\,A\,B_{2}\,,\,A$". The transformer shows an increasing tendency to complete the context with $B_{1}$ as the frequency of token pattern $A\,B_{1}$ in the context grows.
To validate the results derived from the theory, we conducted experiments using real data as well as data generated by the bigram model.
@@
We utilize the analogical reasoning task, one of the most practical tasks for bigram analysis. Specifically, we focus on the "capital-world" analogy type of questions from the Google Analogy Dataset [36], which contains a total of 19,544 questions. This subset includes 232 unique vocabulary items. Denoting a word $A$ (e.g., Tokyo) and its corresponding analogy counterpart $A^{*}$ (e.g., Japan), the prompts were constructed in the form $A\,A^{*}\,,\,B\,B^{*}\,,\,C\,C^{*}\,,\dots$, with pairs separated by commas. Using these generated prompts, we trained $\operatorname{TF}_{\text{rpe}}$. For details on the training and dataset setup, please refer to Appendix D.
@@
As shown in Corollary 1, the model predicts either a token based on global knowledge or the token that appears in the context. To confirm this, we randomly sampled $A$ from the capitals in the Google Analogy Dataset and constructed prompts using word pairs $A\,B_{1}$ and $A\,B_{2}$ that were not learned as global knowledge. The resulting prompts were of the form $A\,B_{1}\,,\dots,\,A\,B_{1}\,,\,A\,B_{2}\,,\dots,\,A\,B_{2}\,,\,A$, which were then used for predicting the next token. We generated 1,000 prompts and calculated the proportion of cases where either $B_{1}$ or $B_{2}$ was predicted, while controlling the frequency of $B_{1}$ and $B_{2}$ in the context.
The result is shown in Fig. 3. As the number of token pattern $A\,B_{1}$ increases, the model predicts $B_{1}$ as the next token more frequently. Furthermore, it can be observed that the prediction trend reverses when the frequency of $B_{1}$ and $B_{2}$ in the context is nearly equal. Additionally, we observe from global ratio that the model predicted some vocabulary $B$ from global knowledge. Especially, the figure illustrates that it is more likely to output global knowledge when the number of $B_{1}$ and $B_{2}$ are the same, or when the number of $B_{1}$ or $B_{2}$ is almost maximum. Please refer to Appendix E.2 for further discussion on these phenomena.
@@ 6 Conclusion
We analyzed the influence of interaction between global bigram and in-context knowledge to the two-layer transformer through the lens of associative memory. We theoretically and empirically verified that relative positional encoding enables transformers to generalize longer sequences, and that how the next token prediction is conducted within the transformer that was trained to store global knowledge and to have induction head.
@@ 7 Limitations
Second, while our focus on induction heads provides valuable insights into in-context learning ICL, it is important to note that ICL leverages additional computational circuits beyond induction heads. As such, our explanation captures only part of the broader mechanisms underlying ICL, leaving room for further investigation into other contributing factors.
@@ Appendix A Additional related work
Appendix B Learning of ideal transformer
B.1 Notation table
@@ B.2 Setup
B.3 Theoretical analysis of learning
@@
Under the setup described in Sec. B.2, a two-layer attention only transformer with relative positional encoding learns associations of the associative memory transformer, i.e., the memories $W_{O}^{2},W_{K}^{2}$ and $W_{K}^{1}$ can be written as the sum of the terms that constitute the associative memory transformer.
@@
Combining eq. 17 and B.3, the near-orthogonality of Gaussian vectors gives us
@@
Since the training of $W_{O}^{2}$ is finished, we have Eq. B.3. When $W_{K}^{1}=W_{K}^{2}=O_{d\times d}$, Lemma 4 demonstrates that the next token prediction is distributed among all tokens, which means that $\hat{p}(v\mid x)=1/V$ for any vocabulary $v$. According to Lemma 5, one step of gradient descent yields
@@
where $\bar{x}^{(1,0)}=\frac{1}{T}\sum_{t=1}^{T}x_{t}^{(1,0)}$. Using the formula 22, we obtain
@@
Similarly, we have $\Delta_{k,j}^{q}=O\left(\frac{1}{N}\right)$ and $\Delta_{k,j}^{r}=O\left(\frac{T}{N}\right)$ based on the discussion in Bietti et al. [4]. By combining these analysis with Eq. B.3, we multiply $(\Phi_{1}w_{E}(i))^{\top}$ from left and $w_{E}(j)$ from right to Eq. 24 and we establish that
@@
Appendix C Other theoretical proofs
@@
Consider a two-layer attention only transformer that meets the initialization condition in Sec. B.2, and the input sequence of length $T$ ends with the trigger token $z_{T}=q$ at its second occurrence. Even after the gradient descent step of $W_{O}^{2}$ and $W_{K}^{2}$, the prediction $\hat{p}(v\mid z_{1:T})=1/V$ for any $v\in\mathcal{V}$.
@@
Consider a two-layer attention only transformer that meets the initialization condition in Sec. B.2, and the input sequence of length $T$ ends with the trigger token $z_{T}=q$ at its second occurrence. After the gradient descent step of $W_{O}^{2}$, the prediction $\hat{p}(v\mid z_{1:T})=1/V$ for any $v\in\mathcal{V}$.
@@ C.1 Supporting lemmas
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. For any $t\in[1,T]$, and any vocabulary $k\in\mathcal{V}$, we have
@@
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. Fix some $t\in[1,T]\subset\mathbb{N}$. For any $s\in[1,t-2]\cup\{t\}$, and vocabulary $k\in\mathcal{V}$, we have
@@
Combining Eq. C.1 and 37, we obtain the result. ∎
@@
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. For any vocabulary $k\in\mathcal{V}$, we have
@@
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. Fix some $t\in[1,T]\subset\mathbb{N}$. For any $s\in[1,t-2]\cup\{t\}$, and vocabulary $k\in\mathcal{V}$,
@@
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. For any vocabulary $k\in\mathcal{V}$, and $t\in[1,T]\subset\mathbb{N}$, we have
@@
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. For any vocabulary $k\in\mathcal{V}$, and $t\in[1,T]\subset\mathbb{N}$, we have
@@
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. For any vocabulary $k\in\mathcal{V}$, and $t\in[1,T]\subset\mathbb{N}$, we have
@@
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. For any vocabulary $k\in\mathcal{V}$, and $t\in[1,T]\subset\mathbb{N}$, we have
@@
We already have the result as to $x_{t_{q}}$ in Eq. C.1:
@@
For this scenario, we have seen that $t_{q}=t=T-1$ is the only possible way for the expectation to take non-zero value. For $x_{t_{q}}$, we can observe Eq. C.1 and obtain
@@
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. Then, we have
@@
Given a sequence $z_{1:T}$ that satisfies Sec. B.2. Then, we have
@@
Suppose a two-layer transformer is a stronger associative memory transformer as in Eqs. 10, 11, and 12. Given a length-$T$ sequence $z_{1:T}$ where the last token is $z_{T}=q$, let $f(v)$ be the number of token pattern "$qv$" appearing in the sequence for vocabulary $v\in\mathcal{V}$. For large enough $\tau_{1}$ and $\tau_{2}$, the logits $\xi_{v}$ can be expressed as follows:
@@
C.3 Induction head without positional encoding

### v2

@@ 4.1 Oversight of patterns
For vocabulary $v\neq v_{1},v_{2}$, the logit $\xi_{v}$ is determined only by $\log{\pi_{b}(v\mid q)}$ and this means that the candidate for the next token is $\operatorname*{arg\,max}_{v\in\mathcal{V}\setminus\{v_{1},v_{2}\}}\pi_{b}(v\mid q)$. On the other hand, The logits for $v_{1}$ and $v_{2}$ have the other term in Prop. 3, which makes them other candidates for the next token. Overall, Corollary 1 states that the model output depends on the bigram conditionals $\pi_{b}(\cdot\mid q)$ and the frequency of token patterns $qv_{1}$ and $qv_{2}$ in the context.
@@ 5 Experiments
This section empirically examines how positional encoding affects the transformer’s ability to capture patterns in input sequences, as discussed in Thm. 1 and 2, and how the transformer’s output changes depending on the proportion of pattern occurrences within the context, as stated in Prop. 3. Henceforth, we denote by $\operatorname{TF}_{\text{ape}}$ and $\operatorname{TF}_{\text{rpe}}$ two-layer transformer with APE, RPE.
@@ 5.1 Oversight of in-context knowledge
(a) Even when trained with sequences of length $256$, the attention block in the first layer of $\operatorname{TF}_{\text{APE}}$ fails to attend to the previous token at positions $t>128$, while $\operatorname{TF}_{\text{RPE}}$ attends to previous tokens regardless of the positions.
(b) Ratio of predicting $B_{1}$ and $B_{2}$ given the prompt "$A\,B_{1}\,,\dots,\,A\,B_{1}\,,\,A\,B_{2}\,,\dots,\,A\,B_{2}\,,\,A$". The transformer shows an increasing tendency to complete the context with $B_{1}$ as the frequency of token pattern $A\,B_{1}$ in the context grows.
Figure 1: (a) Comparison of a two-layer transformer with APE (Thm. 1) and RPE (Thm. 2) capturing previous token information. (b) Output behavior as the frequency of the token pattern $A\,B_{1}$ increases in the input sequence (Prop. 3).
@@
We investigated whether the training of transformers directs attention towards the previous token in the first attention layer. Specifically, $\operatorname{TF}_{\text{rpe}}$ and $\operatorname{TF}_{\text{ape}}$ were trained using sequences of length $256$ generated according to the sequence generation rule. Details regarding the hyperparameters and sequence generation process are provided in Appendix E.
For evaluation of the learned associative memory $\sum_{(i,j)\in M}u_{i}v_{j}^{\top}$, we employed the memory recall metric [13, 20, 5]:
@@
For $\operatorname{TF}_{\text{rpe}}$, we want each vocabulary $w_{E}(k)$ to be paired with the information of the previous position $r_{-1}$. Hence, $R(W_{K}^{1})$ is computed as
$\frac{1}{\|Q\|}\sum_{k\in Q}\mathbf{1}\{\operatorname*{arg\,max}_{i^{\prime}}r_{i^{\prime}}^{\top}W_{K}^{1}w_{E}(k)=-1\}.$
Similarly, we have for $\operatorname{TF}_{\text{ape}}$
@@
where $T$ is the maximum input length for the transformer.
As we can see from Fig. 1(a), over $90\%$ of desired associations are stored in $W_{K}^{1}$ for $\operatorname{TF}_{\text{rpe}}$, while $\operatorname{TF}_{\text{ape}}$ fails to remember most of the associations at positions $256128<t<$ 256 even though the model was trained with sequences of length $256$.
@@ Discussion
To validate the results derived from the theory, we conducted experiments using real data.
@@
We utilize the analogical reasoning task, one of the most practical tasks for bigram analysis. Specifically, we focus on the "capital-world" analogy type of questions from the Google Analogy Dataset [38], which contains a total of 19,544 questions. This subset includes 232 unique vocabulary items. Denoting a word $A$ (e.g., Tokyo) and its corresponding analogy counterpart $A^{*}$ (e.g., Japan), the prompts were constructed in the form $A\,A^{*}\,,\,B\,B^{*}\,,\,C\,C^{*}\,,\dots$, with pairs separated by commas. Using these generated prompts, we trained $\operatorname{TF}_{\text{rpe}}$. For details on the training and dataset setup, please refer to Appendix E.
@@
As shown in Corollary 1, the model predicts either a token based on global knowledge or the token that appears in the context. To confirm this, we randomly sampled $A$ from the capitals in the Google Analogy Dataset and constructed prompts using word pairs $A\,B_{1}$ and $A\,B_{2}$ that were not learned as global knowledge. The resulting prompts were of the form $A\,B_{1}\,,\dots,\,A\,B_{1}\,,\,A\,B_{2}\,,\dots,\,A\,B_{2}\,,\,A$, which were then used for predicting the next token. We generated 1,000 prompts and calculated the proportion of cases where either $B_{1}$ or $B_{2}$ was predicted, while controlling the frequency of $A\,B_{1}$ and $A\,B_{2}$ in the context.
The result is shown in Fig. 1(b). As the number of token pattern $A\,B_{1}$ increases, the model predicts $B_{1}$ as the next token more frequently. Furthermore, it can be observed that the prediction trend reverses when the frequency of $B_{1}$ and $B_{2}$ in the context is nearly equal. Additionally, we observe from global ratio that the model predicted some vocabulary $B$ from global knowledge. Especially, the figure illustrates that it is more likely to output global knowledge when the number of $B_{1}$ and $B_{2}$ are the same, or when the number of $B_{1}$ or $B_{2}$ is almost maximum. Please refer to Appendix F.2 for further discussion on these phenomena.
@@ 6 Conclusion
We analyzed the influence of interaction between global bigram and in-context knowledge to the two-layer transformer through the lens of associative memory. We theoretically and empirically verified that relative positional encoding empowers transformers with the ability to capture information in longer sequences, and that how the next token prediction is conducted within the transformer that was trained to store global knowledge and to have induction head.
@@ 7 Limitations
Second, while our focus on induction heads provides valuable insights into ICL, it is important to note that ICL leverages additional computational circuits beyond induction heads. As such, our explanation captures only part of the broader mechanisms underlying ICL, leaving room for further investigation into other contributing factors.
@@ Appendix A Additional related work
Appendix B MLP as a key-value memory
Geva et al. [19] discovered that the feed-forward layers in transformer-based language models function as key-value memory. In a transformer, the feed-forward layer is typically computed as
$F(x)=W_{2}(\operatorname{ReLU(W_{1}x)}),$
where $x\in\mathbb{R}^{d}$ is the input vector, and $W_{1}\in\mathbb{R}^{d_{h}\times d},W_{2}\in\mathbb{R}^{d_{o}\times d_{h}}$. According to Geva et al. [19], the rows of $W_{1}$ (keys) serve to detect specific patterns in the input sequence, while the columns of $W_{2}$ (values) have been shown to represent distributions over the output vocabulary.
Bietti et al. [5] uses a transformer architecture that uses a linear projection $W_{F}$ instead of an MLP layer.
$W_{F}=\sum_{v=1}^{V}\sum_{u=1}^{V}\log{\pi_{b}(u\mid v)}w_{U}(u)w_{E}(v)^{\top}.$
Due to the near-orthogonality of the embedding vectors, this projection maps an input $x=w_{E}(v)$ to the output $\sum_{u=1}^{V}\log{\pi_{b}(u\mid v)}w_{U}(u)$. This enables the matrix $W_{F}$ to function as an associative memory that stores the pairs of an embedding vector and corresponding global knowledge.
Now, suppose we construct the matrices $W_{1}$ and $W_{2}$ as follows:
$W_{1}=\begin{pmatrix}w_{E}(v_{1})^{\top}\\ w_{E}(v_{2})^{\top}\\ \vdots\\ w_{E}(v_{V})^{\top}\end{pmatrix},$ (11)
$W_{2}=\begin{pmatrix}\sum_{u=1}^{V}\log{\pi_{b}(u\mid v_{1})}w_{U}(u)^{\top}\\ \quad\sum_{u=1}^{V}\log{\pi_{b}(u\mid v_{2})}w_{U}(u)^{\top}\\ \vdots\\ \sum_{u=1}^{V}\log{\pi_{b}(u\mid v_{V})}w_{U}(u)^{\top}\end{pmatrix}^{\top},$ (12)
Each row of $W_{1}$ acts as a detector for an input pattern $w_{E}(v)$, such that only the corresponding entry is activated when $w_{E}(v_{i})$ is provided as input. Each row of $W_{2}$, in turn, encodes global knowledge, i.e., the distribution $\pi_{b}$ over the output vocabulary conditioned on the corresponding input token.
We can easily find that the resulting computation matches exactly that of the linear projection $W_{F}$ under the near-orthgonality condition.
Therefore, this construction confirms that the feed-forward layer operates as a key-value memory.
Appendix C Learning of ideal transformer
C.1 Notation table
@@ C.2 Setup
C.3 Theoretical analysis of learning
@@
Under the setup described in Sec. C.2, a two-layer attention only transformer with relative positional encoding learns associations of the associative memory transformer, i.e., the memories $W_{O}^{2},W_{K}^{2}$ and $W_{K}^{1}$ can be written as the sum of the terms that constitute the associative memory transformer.
@@
Combining eq. 16 and C.3, the near-orthogonality of Gaussian vectors gives us
@@
Since the training of $W_{O}^{2}$ is finished, we have Eq. C.3. When $W_{K}^{1}=W_{K}^{2}=O_{d\times d}$, Lemma 4 demonstrates that the next token prediction is distributed among all tokens, which means that $\hat{p}(v\mid x)=1/V$ for any vocabulary $v$. According to Lemma 5, one step of gradient descent yields
@@
where $\bar{x}^{(1,0)}=\frac{1}{T}\sum_{t=1}^{T}x_{t}^{(1,0)}$. Using the formula 21, we obtain
@@
Similarly, we have $\Delta_{k,j}^{q}=O\left(\frac{1}{N}\right)$ and $\Delta_{k,j}^{r}=O\left(\frac{T}{N}\right)$ based on the discussion in Bietti et al. [5]. By combining these analysis with Eq. C.3, we multiply $(\Phi_{1}w_{E}(i))^{\top}$ from left and $w_{E}(j)$ from right to Eq. 23 and we establish that
@@
Appendix D Other theoretical proofs
@@
Consider a two-layer attention only transformer that meets the initialization condition in Sec. C.2, and the input sequence of length $T$ ends with the trigger token $z_{T}=q$ at its second occurrence. Even after the gradient descent step of $W_{O}^{2}$ and $W_{K}^{2}$, the prediction $\hat{p}(v\mid z_{1:T})=1/V$ for any $v\in\mathcal{V}$.
@@
Consider a two-layer attention only transformer that meets the initialization condition in Sec. C.2, and the input sequence of length $T$ ends with the trigger token $z_{T}=q$ at its second occurrence. After the gradient descent step of $W_{O}^{2}$, the prediction $\hat{p}(v\mid z_{1:T})=1/V$ for any $v\in\mathcal{V}$.
@@ D.1 Supporting lemmas
Given a sequence $z_{1:T}$ that satisfies Sec. C.2. For any $t\in[1,T]$, and any vocabulary $k\in\mathcal{V}$, we have
@@
Given a sequence $z_{1:T}$ that satisfies Sec. C.2. Fix some $t\in[1,T]\subset\mathbb{N}$. For any $s\in[1,t-2]\cup\{t\}$, and vocabulary $k\in\mathcal{V}$, we have
@@
Combining Eq. D.1 and 36, we obtain the result. ∎
@@
Given a sequence $z_{1:T}$ that satisfies Sec. C.2. For any vocabulary $k\in\mathcal{V}$, we have
@@
Given a sequence $z_{1:T}$ that satisfies Sec. C.2. Fix some $t\in[1,T]\subset\mathbb{N}$. For any $s\in[1,t-2]\cup\{t\}$, and vocabulary $k\in\mathcal{V}$,
@@
Given a sequence $z_{1:T}$ that satisfies Sec. C.2. For any vocabulary $k\in\mathcal{V}$, and $t\in[1,T]\subset\mathbb{N}$, we have
@@
Given a sequence $z_{1:T}$ that satisfies Sec. C.2. For any vocabulary $k\in\mathcal{V}$, and $t\in[1,T]\subset\mathbb{N}$, we have
@@
Given a sequence $z_{1:T}$ that satisfies Sec. C.2. For any vocabulary $k\in\mathcal{V}$, and $t\in[1,T]\subset\mathbb{N}$, we have
@@
Given a sequence $z_{1:T}$ that satisfies Sec. C.2. For any vocabulary $k\in\mathcal{V}$, and $t\in[1,T]\subset\mathbb{N}$, we have
@@
We already have the result as to $x_{t_{q}}$ in Eq. D.1:
@@
For this scenario, we have seen that $t_{q}=t=T-1$ is the only possible way for the expectation to take non-zero value. For $x_{t_{q}}$, we can observe Eq. D.1 and obtain
@@
Given a sequence $z_{1:T}$ that satisfies Sec. C.2. Then, we have
@@
Given a sequence $z_{1:T}$ that satisfies Sec. C.2. Then, we have
@@
Suppose a two-layer transformer is a stronger associative memory transformer as in Def. 3. Given a length-$T$ sequence $z_{1:T}$ where the last token is $z_{T}=q$, let $f(v)$ be the number of token pattern "$qv$" appearing in the sequence for vocabulary $v\in\mathcal{V}$. For large enough $\tau_{1}$ and $\tau_{2}$, the logits $\xi_{v}$ can be expressed as follows:
@@
D.3 Induction head without positional encoding